how would find the value of 3[f(-2)] if f(x)=4x+2?

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

find the value of 3[f(-2)] if f(x)=4x+2?
--------------
f(x)=4x+2
f(-2) = 4*-2+2 = -8+2 = -6
---
3 = 3*-6 = -18
===========================
Cheers,
Stan H.
============

Answer by solver91311(24713)



Replace in the definition of with -2, then calculate the value of . Multiply the result by 3.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism